# CP Violation in $D\to KK$ Decays: A Comparative Analysis of Triplet and Sextet Diquarks

### v1

@@ Abstract
Recent measurements of the CP asymmetry in the decay $D^{0}\rightarrow K_{S}^{0}K_{S}^{0}$ by the CMS collaboration, $A_{CP}(K_{S}^{0}K_{S}^{0})=(6.2\pm 3.0\pm 0.2\pm 0.8)\%$, and by LHCb, $A_{CP}(D^{0}\to K_{S}^{0}K_{S}^{0})=(1.86\pm 0.23\pm 0.11)\%$, suggest possible deviations from Standard Model (SM) expectations, which predict asymmetries below the percent level. This singly Cabibbo-suppressed decay is particularly sensitive to new physics, as the leading amplitudes vanish in the exact U-spin symmetry limit and the process is dominated by W-exchange topologies. We investigate scalar diquark contributions to this decay, comparing color-sextet and color-triplet representations. We find that the color-sextet diquark, characterized by a symmetric color structure $(C_{1}^{\mathrm{NP}}=C_{2}^{\mathrm{NP}})$, avoids color suppression and can generate CP asymmetries in the range $0.5\%$$1.5\%$ for a diquark mass of order 1 TeV. In contrast, the color-triplet contribution is strongly suppressed due to destructive interference from its antisymmetric color structure. We further show that a flavor hierarchy in the sextet couplings, with $\lambda_{ud}>\lambda_{us}$, can simultaneously account for the observed deviation from the U-spin sum rule in $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$ and the measured CP asymmetry in $D^{0}\to K_{S}^{0}K_{S}^{0}$. These results identify color-sextet scalar diquarks as viable candidates for explaining enhanced CP violation in charm decays.
@@ I Introduction
The experimental study of CP violation in charm decays has recently entered a new era of precision, driven by the large datasets collected by the LHCb and CMS experiments. Of particular interest is the decay channel $D^{0}\to K_{S}^{0}K_{S}^{0}$, which provides a uniquely sensitive probe of new sources of CP violation. Recent measurements by the LHCb collaboration have reported a direct CP asymmetry [1]:
$A_{CP}(D^{0}\to K_{S}^{0}K_{S}^{0})=(1.86\pm 0.23\pm 0.11)\%,$
@@
Theoretical interest in charm CP violation has been further stimulated by recent experimental indications of deviations from U-spin sum rules relating the CP asymmetries of the SCS decay modes $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$. Global fits to experimental data suggest a violation of these relations at the level of approximately $2.7\sigma$, with both asymmetries observed to be positive, contrary to the leading-order SM expectation [6]. These observations point toward the intriguing possibility of new physics contributions to charm decay amplitudes.
@@ IV Phenomenology of Scalar Diquarks in $D^{0}\to K_{S}^{0}K_{S}^{0}$
Experimentally, LHCb has measured
$A_{CP}(D^{0}\to K_{S}^{0}K_{S}^{0})=(1.86\pm 0.23\pm 0.11)\%,$ (18)

### v2

# arXiv:2602.21181v2[hep-ph] 02 Mar 2026

@@ Abstract
Recent measurements of the CP asymmetry in the decay $D^{0}\rightarrow K_{S}^{0}K_{S}^{0}$ by the CMS collaboration, $A_{CP}(K_{S}^{0}K_{S}^{0})=(6.2\pm 3.0\pm 0.2\pm 0.8)\%$, and by LHCb, $A_{CP}(D^{0}\to K_{S}^{0}K_{S}^{0})=(1.86\pm 1.04\pm 0.41)\%$, suggest possible deviations from Standard Model (SM) expectations, which predict asymmetries below the percent level. This singly Cabibbo-suppressed decay is particularly sensitive to new physics, as the leading amplitudes vanish in the exact U-spin symmetry limit and the process is dominated by W-exchange topologies. We investigate scalar diquark contributions to this decay, comparing color-sextet and color-triplet representations. We find that the color-sextet diquark, characterized by a symmetric color structure $(C_{1}^{\mathrm{NP}}=C_{2}^{\mathrm{NP}})$, avoids color suppression and can generate CP asymmetries in the range $0.5\%$$1.5\%$ for a diquark mass of order 1 TeV. In contrast, the color-triplet contribution is strongly suppressed due to destructive interference from its antisymmetric color structure. We further show that a flavor hierarchy in the sextet couplings, with $\lambda_{ud}>\lambda_{us}$, can simultaneously account for the observed deviation from the U-spin sum rule in $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$ and the measured CP asymmetry in $D^{0}\to K_{S}^{0}K_{S}^{0}$. These results identify color-sextet scalar diquarks as viable candidates for explaining enhanced CP violation in charm decays.
@@ I Introduction
The experimental study of CP violation in charm decays has recently entered a new era of precision, driven by the large datasets collected by the LHCb and CMS experiments. Of particular interest is the decay channel $D^{0}\to K_{S}^{0}K_{S}^{0}$, which provides a uniquely sensitive probe of new sources of CP violation. Recent measurements by the LHCb collaboration have reported a direct CP asymmetry [1, 2]:
$A_{CP}(D^{0}\to K_{S}^{0}K_{S}^{0})=(1.86\pm 1.04\pm 0.41)\%,$
@@
Theoretical interest in charm CP violation has been further stimulated by recent experimental indications of deviations from U-spin sum rules relating the CP asymmetries of the SCS decay modes $D^{0}\to K^{+}K^{-}$ and $D^{0}\to\pi^{+}\pi^{-}$. Global fits to experimental data suggest a violation of these relations at the level of approximately $2.7\sigma$, with both asymmetries observed to be positive, contrary to the leading-order SM expectation [7, 8, 9]. These observations point toward the intriguing possibility of new physics contributions to charm decay amplitudes.
@@ IV Phenomenology of Scalar Diquarks in $D^{0}\to K_{S}^{0}K_{S}^{0}$
Experimentally, LHCb has measured [1]:
$A_{CP}(D^{0}\to K_{S}^{0}K_{S}^{0})=(1.86\pm 1.04\pm 0.41)\%,$ (18)